$SPAD/src/input draw.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
)clear all

-- Two dimensional function plots
draw(sin(11*x),x = 0..2*%pi)
draw(sin(11*x),x = 0..2*%pi,adaptive == false,title == "Non-adaptive plot")
draw(sin(11*x),x = 0..2*%pi,toScale == true,title == "Drawn to scale")
draw(sin(11*x),x = 0..2*%pi,coordinates == polar,title == "Polar plot")
draw(tan x,x = -6..6,title == "Clipping on")
draw(tan x,x = -6..6,clip == false,title == "Clipping off")
-- Parametric plane curves
draw(curve(9*sin(3*t/4),8*sin(t)),t = -4*%pi..4*%pi,_
           toScale == true,title == "Lissajous curve")
draw(curve(sin(5*t),t),t = 0..2*%pi,title == "Parametric curve")
draw(curve(sin(5*t),t),t = 0..2*%pi,_
           coordinates == polar,title == "Parametric polar curve")
draw(curve(t,2),t = -3..3,coordinates == parabolic,title == "Parabolic plot")
-- Parametric space curves
draw(curve(sin(t)*cos(3*t/5),cos(t)*cos(3*t/5),cos(t)*sin(3*t/5)),_
           t = 0..15*%pi,title == "Parametric curve")
draw(curve(sin(t)*cos(3*t/5),cos(t)*cos(3*t/5),cos(t)*sin(3*t/5)),_
           t = 0..15*%pi,tubeRadius == .15,title == "Tube around curve")
draw(curve(4,t,t),t = -9..9,coordinates == cylindrical,title == "Spiral")
draw(curve(1,t/7,t/5),t = -70*%pi..70*%pi,coordinates == spherical,_
           title == "Gnarly curve on sphere")
-- Three dimensional function plots
draw(cos(x*y),x = -3..3,y = -3..3)
colorFunction1(x:DFLOAT,y:DFLOAT):DFLOAT == x
draw(cos(x*y),x = -3..3,y = -3..3,colorFunction == colorFunction1)

colorFunction2(x:DFLOAT,y:DFLOAT):DFLOAT == x**2 + y**2
draw(cos(x*y),x = -3..3,y = -3..3,colorFunction == colorFunction2)
colorFunction3(x:DFLOAT,y:DFLOAT,z:DFLOAT):DFLOAT == -z
draw(cos(x*y),x = -3..3,y = -3..3,colorFunction == colorFunction3)
draw(1,u = 0..2*%pi,v = 0..%pi,coordinates == spherical,title == "Sphere")
draw(1,u = 0..2*%pi,v = 0..%pi,coordinates == spherical,_
             colorFunction == colorFunction2,title == "Sphere")
draw(1,u = -%pi/2..%pi/2,v = 0..2*%pi,coordinates == oblateSpheroidal(1$DFLOAT),_
           title == "Oblate spheroidal coordinates")
-- Parametric surfaces
draw(surface(5*sin(u)*cos(v),4*sin(u)*sin(v),3*cos(u)),u = 0..%pi,v = 0..2*%pi,_
           title == "Which came first?")

———-

References^{[1] nothing}

5